The difference in 2 numbers is 92. When the larger is subtracted from 4 times the smaller, the result is 151. Find the numbers. Define the variables. Write 2 equations and solve.
Answer by JulietG(1812)

A - B = 92
4B - A = 151
In the second equation, add A to each side
4B = A + 151
Subtract 151 from each side
4B - 151 = A
Now plug that value into the top equation
(4B-151) - B = 92
Subtract the B
3B - 151 = 92
Add 151 to each side
3B = 243
Divide each side by 3
B = 81
If B is 81, then A must B 92+81, or 173